### v1

# FedMLAC: Mutual Learning Driven Heterogeneous Federated Audio Classification

Abstract
Federated Learning (FL) provides a privacy-preserving paradigm for training audio classification (AC) models across distributed clients without sharing raw data. However, Federated Audio Classification (FedAC) faces three critical challenges that substantially hinder performance: data heterogeneity, model heterogeneity, and data poisoning. While prior works have attempted to address these issues, they are typically treated independently, lacking a unified and robust solution suited to real-world federated audio scenarios. To bridge this gap, we propose FedMLAC, a unified mutual learning framework designed to simultaneously tackle these challenges in FedAC. Specifically, FedMLAC introduces a dual-model architecture on each client, comprising a personalized local AC model and a lightweight, globally shared Plug-in model. Through bidirectional knowledge distillation, the Plug-in model enables global knowledge transfer while adapting to client-specific data distributions, thus supporting both generalization and personalization. To further enhance robustness against corrupted audio data, we develop a Layer-wise Pruning Aggregation (LPA) strategy that filters unreliable Plug-in model updates based on parameter deviations during server-side aggregation. Extensive experiments on four diverse audio classification benchmarks—spanning both speech and non-speech tasks—demonstrate that FedMLAC consistently outperforms existing state-of-the-art methods in terms of classification accuracy and robustness to noisy data.
keywords:
Federated Learning , Audio Classification , Mutual Learning , Data Heterogeneity , Model Heterogeneity , Data Poisoning.
[…]
[0]organization=School of Computer Science, McGill University,city=Montreal, country=Canada
[…]
[1]organization=Mila - Quebec AI Institute,city=Montreal, country=Canada
[…]
[2]organization=School of Mathematics, Physics and Computing, University of Southern Queensland,city=Springfield, country=Australia
[…]
[3]organization=School of Mathematics, Physics and Computing, University of Southern Queensland,city=Toowoomba, country=Australia
[…]
[4]organization=Data61, CSIRO,city=Melbourne, country=Australia
1 Introduction
Audio classification (AC), defined as the task of assigning semantic labels to audio signals, has become a foundational component of modern intelligent systems [1]. It supports a wide range of applications, including speech command recognition [2], emotion detection [3], environmental sound classification, and bioacoustic monitoring [4]. These tasks span diverse acoustic domains and require models to capture fine-grained temporal and spectral features from highly variable and often noisy data [5, 6]. Advances in deep learning—particularly convolutional neural networks (CNNs), recurrent neural networks (RNNs), and transformer-based architectures—have significantly improved classification performance under both supervised and weakly supervised settings [1], relying on large-scale, diverse datasets to learn robust acoustic representations from raw waveforms or spectrograms. In practice, however, such datasets are increasingly generated and retained locally on heterogeneous edge devices such as smartphones, wearables, smart speakers, and IoT sensors [7], rather than being centrally collected. This shift toward decentralized data generation poses both technical and ethical challenges to centralized training, especially in light of privacy concerns, communication overhead, and legal regulations such as the GDPR [8] and the EU AI Act [9]. These constraints underscore the need for alternative learning frameworks that can operate effectively in decentralized environments while ensuring data privacy and communication efficiency.
One of the most promising solutions to address the challenges of decentralized and privacy-sensitive audio data is Federated Learning (FL), a distributed training paradigm that enables multiple clients to collaboratively learn a shared model without exchanging raw data [10]. By performing computation locally and aggregating only model updates, FL reduces communication overhead and mitigates privacy risks, making it particularly attractive for deployment on edge devices such as smartphones, wearables, and IoT sensors [11, 12]. This paradigm aligns well with the characteristics of modern audio data ecosystems, where large volumes of diverse audio signals remain distributed across heterogeneous sources [6]. Compared to centralized learning approaches, FL offers inherent advantages in terms of privacy preservation, regulatory compliance, and scalability [13]. These properties make it a natural candidate for large-scale AC tasks under real-world constraints.
In this context, Federated Audio Classification (FedAC) refers to the application of FL to collaboratively train AC models across distributed clients while keeping local audio data private [14, 6, 15, 16, 17]. FedAC enables learning from diverse, sensitive, and geographically dispersed audio sources without requiring centralized access, which is particularly valuable in domains such as healthcare [18], smart environments [16], and personalized voice interfaces [19]. However, applying FL to AC introduces several unique challenges. First, data heterogeneity arises from variations in speaker identity, dialect, background noise, and recording conditions, resulting in highly non-IID data distributions across clients [20, 21, 22]. Second, model heterogeneity stems from differences in device capabilities, runtime environments, and model architectures, complicating coordination and degrading global convergence [23, 24]. Third, data poisoning in the federated setting may arise from both malicious clients and naturally occurring issues such as label errors or input noise, potentially leading to degraded global model performance and unstable convergence [25, 26]. These challenges collectively limit the performance, robustness, and scalability of conventional FL approaches when applied to real-world AC tasks.
Despite growing interest in FL for both speech and non-speech audio applications, existing research tends to address the challenges of data heterogeneity, model heterogeneity, and data poisoning in partial or disjointed ways. A substantial body of work focuses on mitigating non-IID data distributions through strategies such as regularization techniques [27, 28], client selection [29, 30, 31, 32, 33, 34], representation learning [35], or robust aggregation mechanisms [36, 20, 21, 37]. While these techniques improve global model utility under skewed data, they often assume a homogeneous model architecture shared across all clients, which is impractical in real-world deployments involving heterogeneous edge devices. To address system-level diversity, recent studies have explored flexible model configurations, including personalized layers [38, 39, 40], decoupled training [41, 27, 42, 43, 16], incremental learning [44, 45, 46], or sparsified aggregation [47, 48]. However, such methods rarely account for the interaction between model and data heterogeneity, which can cause unstable convergence, degraded generalization, or fairness concerns when local models diverge significantly.
Another important challenge in FedAC is the presence of corrupted or unreliable local data, commonly referred to as data poisoning. Such corruption can arise not only from adversarial clients but also from natural sources, including label inconsistencies in subjective tasks (e.g., emotion recognition) and distorted audio recordings in uncontrolled environments. These issues can significantly impair model stability and convergence, especially when coupled with non-IID data or heterogeneous client models. To mitigate these effects, researchers have proposed defense strategies such as anomaly detection [27, 49, 35, 50], adversarial learning [51], client clustering [52], and model pruning [53, 54]. While effective in isolation, these methods are typically reactive and rarely consider the joint effects of data and model heterogeneity. Moreover, many introduce a trade-off between robustness and utility, often degrading performance on clean or low-resource clients. Thus, achieving resilience to data corruption without sacrificing personalization or fairness remains an open problem in FedAC.
Table 1: Systematic comparison of existing FedAC methods with respect to three core challenges: data heterogeneity, model heterogeneity, and data poisoning. Note: $\checkmark$ (Addressed); $\times$ (Not addressed).
Studies Data Heterogeneity Model Heterogeneity Data Poisoning
[36, 20, 21, 29, 35, 37, 28, 33, 32] $\checkmark$ $\times$ $\times$
[47, 39, 48, 40, 42, 43] $\times$ $\checkmark$ $\times$
[27, 49, 50, 53, 54, 51, 52] $\checkmark$ $\times$ $\checkmark$
[38, 41, 44, 45, 46] $\checkmark$ $\checkmark$ $\times$
FedMLAC (Ours) $\checkmark$ $\checkmark$ $\checkmark$
As summarized in Table 1, existing research efforts in FedAC remain largely fragmented, with most methods targeting only one or two specific challenges. While notable progress has been made in mitigating data heterogeneity or accommodating model diversity, few approaches recognize their interdependence, and even fewer address the added complexity introduced by corrupted or unreliable local data. This gap is especially consequential in AC scenarios, where task subjectivity, noisy environments, and hardware diversity are the norm rather than the exception. The interplay between statistical non-IID distributions, architectural heterogeneity, and data poisoning presents a compounded threat to model convergence, generalization, and fairness—making partial solutions insufficient for deployment in dynamic, real-world settings. To the best of our knowledge, no existing work proposes a unified framework capable of jointly addressing all three challenges in a coordinated and scalable manner. This observation underlines a critical need for integrated approaches that go beyond siloed techniques and holistically enhance the robustness and adaptability of FedAC systems.
To address the intertwined challenges of data heterogeneity, model heterogeneity, and data corruption in FedAC, we propose FedMLAC, a Mutual Learning-driven Federated Audio Classification framework. FedMLAC consists of two key components: FedMLAC-Update and FedMLAC-Aggregation. In FedMLAC-Update, each client maintains a personalized local AC model and a lightweight server-distributed Plug-in model. Through mutual learning, the Plug-in model transfers global knowledge to the local model while simultaneously absorbing client-specific information. This bidirectional distillation ensures that clients benefit from global generalization and local personalization without sharing raw data. Meanwhile, in FedMLAC-Aggregation, the updated Plug-in models from clients are aggregated on the server via a robust Layer-wise Pruning Aggregation (LPA) strategy. This approach selectively integrates trustworthy model layers while pruning extreme deviations—caused by non-IID data or adversarial noise—thereby improving the resilience and stability of the global Plug-in model. Together, these two modules enable FedMLAC to achieve scalable, personalized, and communication-efficient learning across heterogeneous audio environments.
A detailed comparison between FedMLAC and representative federated approaches is presented in Table 1, highlighting the framework’s ability to simultaneously tackle statistical data heterogeneity, model diversity, noisy or corrupted data, and personalization—four challenges that prior methods typically address in isolation. FedMLAC unifies these components into a cohesive and scalable framework: it supports flexible, personalized client models via decoupled optimization; enhances local adaptation and global generalization through mutual knowledge distillation; and achieves robust model aggregation via layer-wise pruning to filter out anomalous or adversarial updates. This integrated design allows FedMLAC to maintain strong performance across highly heterogeneous and dynamic AC environments, as verified in our extensive experimental evaluation.
The main contributions of this work are summarized as follows:
  • 1.
    We propose FedMLAC, a unified federated learning framework for audio classification that is capable of simultaneously addressing three fundamental challenges: data heterogeneity, model heterogeneity, and data poisoning. To the best of our knowledge, this is the first framework to jointly solve all three challenges in the federated audio classification setting.
  • 2.
    To effectively tackle data and model heterogeneity, we introduce a mutual learning mechanism based on a lightweight, globally shared Plug-in model. This Plug-in model enables bidirectional knowledge transfer between global and local models, enhancing generalization across clients while supporting flexible, personalized architectures.
  • 3.
    To enhance robustness against data poisoning—particularly from data noise and label errors—we propose a Layer-wise Pruning Aggregation (LPA) strategy. LPA identifies and prunes abnormal Plug-in model updates at the layer level by measuring parameter deviations across clients, effectively filtering out unreliable contributions during aggregation. This mechanism improves the stability and resilience of the global Plug-in model under noisy and inconsistent local data conditions.
  • 4.
    We provide a theoretical convergence analysis of FedMLAC, showing that the global Plug-in model converges to a stationary point at a rate of $\mathcal{O}(1/\sqrt{T})$ under standard non-convex assumptions. The analysis rigorously characterizes the impact of data heterogeneity, stochastic variance, and layer-wise pruning deviations, thereby offering theoretical guarantees for FedMLAC’s robustness and scalability in heterogeneous federated environments.
  • 5.
    We conduct extensive experiments on four benchmark audio classification datasets, encompassing both speech and non-speech tasks. The results demonstrate that FedMLAC consistently outperforms existing state-of-the-art methods in terms of classification accuracy and robustness to noisy data.
The remainder of this paper is organized as follows. Section 2 provides a comprehensive review of related work on FedAC. Section 3 introduces the problem formulation. Section 4 details the core components and technical implementation of the proposed approach. Section 6 presents the experimental setup, results, and analyses conducted across multiple benchmark datasets. A formal convergence analysis is offered in Section 5. Finally, Section 7 summarizes the key findings and discusses potential directions for future research.
2 Related Work
2.1 Audio Classification
Audio classification (AC) involves automatically assigning predefined semantic labels to audio recordings, such as speech commands, music genres, or environmental sounds. Traditional techniques largely relied on handcrafted features like MFCCs and classical classifiers (e.g., SVMs, GMMs) [55, 56]. With the rise of deep learning, architectures such as CNNs and RNNs have shown superior performance by learning hierarchical representations from raw or spectrogram-based audio [57]. Convolutional Recurrent Neural Networks (CRNNs), which combine the spatial feature extraction power of CNNs with the temporal modeling capabilities of RNNs, have become a widely adopted architecture in audio tagging and event detection [58]. CRNNs typically consist of several convolutional layers for capturing time-frequency local patterns, followed by recurrent layers (e.g., bidirectional GRUs or LSTMs) to model long-term temporal dependencies, and fully connected output layers for final prediction [59].
Despite these advancements, large-scale, high-quality audio datasets remain difficult to collect due to the need for manual annotation and concerns around user privacy. Speech and ambient sound recordings often contain personally identifiable or context-sensitive content, posing regulatory and ethical challenges under data protection frameworks like GDPR [8]. Furthermore, real-world datasets are frequently noisy or weakly labeled, especially when sourced from web platforms [60]. These issues highlight the need for classification systems that are robust to imperfect data, and capable of learning from decentralized or privacy-sensitive sources—motivating recent interest in distributed machine learning paradigms.
2.2 Federated Learning
Federated learning (FL) is a decentralized machine learning paradigm that enables multiple clients to collaboratively train a global model without sharing raw data, thereby preserving user privacy [10, 13]. Its privacy-preserving nature has led to widespread adoption in sensitive domains such as healthcare, mobile services, and finance [13]. However, FL is fundamentally constrained by statistical heterogeneity, where non-IID client data distributions impair convergence and model generalization [11]. To alleviate this, the literature has explored techniques such as regularization-based local training [61, 62], robust aggregation mechanisms [63, 64], and personalized learning frameworks [65]. Simultaneously, the threat of data poisoning—where adversarial clients compromise model integrity—has motivated Byzantine-resilient strategies [66] and trust-aware aggregation schemes [13]. Despite this progress, FL research remains predominantly focused on vision and text benchmarks like CIFAR-10, FEMNIST, and GLUE, with relatively limited efforts targeting audio-based tasks [13]. Given the unique temporal and spectral complexities of audio signals and the privacy sensitivities inherent in speech data, extending FL to acoustic learning scenarios presents both urgent demand and technical novelty. Our work addresses this gap by proposing a federated audio classification framework that accounts for these domain-specific challenges.
2.3 Federated Audio Classification
Recent advancements in Federated Audio Classification (FedAC) have progressively addressed the inherent challenges of decentralized learning over acoustic data. Early efforts in this field largely adopted techniques from computer vision and speech recognition, applying them directly to audio domains without explicitly tailoring to their temporal, spectral, or noise characteristics. For instance, works such as [27, 28] focused on mitigating data heterogeneity through regularization-based federated training, while [29, 30, 32] explored client selection or clustering to group clients with similar data distributions. Representation learning strategies have also emerged [35] to align feature spaces across clients, showing promise in noisy acoustic environments.
More recently, system heterogeneity—stemming from the diversity of devices, compute resources, and model capacities—has gained attention. Personalized FL schemes like [38, 39] allow clients to retain unique local layers, while parameter decoupling [41, 27] and sparse aggregation [47, 48] seek to reduce communication and computation loads. Incremental training paradigms [45, 46] further enhance long-term adaptability, especially under shifting client distributions.
Simultaneously, growing attention has turned to the issue of label noise and unreliable audio data. Both adversarial poisoning [51] and benign corruption—e.g., in emotion recognition or field-recorded data—can derail global convergence. In response, anomaly detection [49, 50], model pruning [53, 54], and unsupervised clustering [52] have been proposed. However, most of these strategies remain isolated to specific challenges, often neglecting the complex interplay between noisy data, non-IID distributions, and heterogeneous model architectures. Few studies attempt to jointly address all three, which leads to limited generalization, unstable training, or diminished fairness on underrepresented clients. As such, holistic frameworks that simultaneously account for statistical, structural, and adversarial variations remain critically underexplored in the current FedAC landscape.
3 Prelimineries
3.1 Problem Setup
Federated Audio Classification (FedAC) aims to train a global Audio Classification (AC) model $\omega^{g}$ collaboratively across multiple clients while keeping each client’s speech data decentralized to preserve privacy. Let there be $N$ clients in the FedAC system, indexed by $k\in\{1,2,\ldots,N\}$. Each client $k$ has a local AC dataset $\mathcal{D}_{k}=\{(\mathbf{x}_{k,j},\mathbf{y}_{k,j})\}_{j=1}^{n_{k}}$, where $\mathbf{x}_{k,j}$ represents the input speech features and $\mathbf{y}_{k,j}$ represents the corresponding label. The global objective of FedAC is to minimize the following loss function:
$_{\omega^{g}}^{min}\mathcal{L}(\omega^{g})=\sum_{k=1}^{N}\frac{n_{k}}{n}\mathcal{L}_{k}(\omega^{g}),$ (1)
where $n=\sum_{k=1}^{N}{n_{k}}$, $n_{k}=|D_{k}|$, $\mathcal{L}_{k}(\omega^{k})=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell(\mathbf{\omega^{k}};\mathbf{x}_{k,j},\mathbf{y}_{k,j})$ denotes the average loss of the ${k^{th}}$ client model $\omega^{k}$ over $D_{k}$, and $l(\cdot)$ represents the training loss (e.g., cross-entropy (CE) loss for AC tasks).
Each client updates its own local AC model $\mathbf{\omega}^{k}$ by performing local training on its dataset $\mathcal{D}_{k}$ using Stochastic Gradient Descent (SGD). The local update rule is:
$\omega^{k}_{t+(i+1)}=\omega^{k}_{t+(i)}-{\eta_{t+(i)}}\nabla{\mathcal{L}_{k}}(\omega^{k}_{t+(i)}),i=0,1,...,E-1,$ (2)
where $\eta$ is the learning rate, $\omega_{t}^{k}$ represents the $k^{th}$ client model in the $t^{th}$ communication round. and $E$ denotes the local training epochs.
After local training, each client sends its updated model $\mathbf{\omega}_{t+1}^{k}$ to the central server, which performs aggregation to update the global model. The most commonly used aggregation method is Federated Averaging (FedAvg) [10], defined as:
$\mathbf{\omega}_{t+1}^{g}=\sum_{k=1}^{N}\frac{n_{k}}{n}\mathbf{\omega}_{t+1}^{k}.$ (3)
The basic FedAC framework requires that each client’s AC model $\mathbf{\omega}^{k}$ maintains the same model architecture. Furthermore, the heterogeneity of distributed AC data leads to an inconsistency between the learning objectives of client models and the global FedAC training objective. These challenges significantly reduce the practicality and stability of existing FedAC frameworks.
3.2 Mutual Learning
Mutual Learning is a collaborative training mechanism where two or more models interactively learn from each other by exchanging predictions during training. It enhances the learning capability of individual models by encouraging consistency and improving generalization [67].
For simplicity, consider the case of two models, $\mathbf{\omega}^{i}$ and $\mathbf{\omega}^{j}$, learning from their respective datasets $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$. The mutual learning objective for each model combines a task-specific loss and a consistency loss. The total loss for $\mathbf{\omega}^{i}$ and $\mathbf{\omega}^{j}$ is defined as:
$\mathcal{L}_{i}&=\mathcal{L}_{\text{task}}(\mathbf{\omega}^{i};\mathcal{D}_{i})+\lambda\mathcal{L}_{\text{KL}}(\mathit{p}_{i},\mathit{p}_{j})\\ \mathcal{L}_{j}&=\mathcal{L}_{\text{task}}(\mathbf{\omega}^{j};\mathcal{D}_{j})+\lambda\mathcal{L}_{\text{KL}}(\mathit{p}_{j},\mathit{p}_{i})\end{aligned}\right.,$ (1)
where $\mathcal{L}_{\text{task}}(\mathbf{\omega}^{i};\mathcal{D}_{i})$ is the task-specific loss (e.g., CE loss) for model $\omega^{i}$ on dataset $\mathcal{D}_{i}$. $\mathcal{L}_{\text{KL}}(\mathit{p}_{i},\mathit{p}_{j})$ is the Kullback-Leibler (KL) divergence between the output distributions $\mathit{p}_{i}$ and $\mathit{p}_{j}$ of models $\mathbf{\omega}^{i}$ and $\mathbf{\omega}^{j}$, defined as:
$\mathcal{L}_{\text{KL}}(\mathit{p}_{i},\mathit{p}_{j})=\sum_{c}\mathit{p}_{i}(c)\log\frac{\mathit{p}_{i}(c)}{\mathit{p}_{j}(c)},$ (4)
where $\mathit{p}_{i}(c)$ represents the probability of class $c$ predicted by model $\mathbf{\omega}^{i}$. $\lambda$ is a hyperparameter balancing the task loss and consistency loss.
The models $\mathbf{\omega}^{i}$ and $\mathbf{\omega}^{j}$ are trained iteratively to minimize their respective losses $\mathcal{L}_{i}$ and $\mathcal{L}_{j}$, leading to mutual enhancement of their predictive performance. This learning paradigm is particularly beneficial in scenarios with model heterogeneity.
4 Methodology
4.1 FedMLAC Overview
Our proposed method consists of two main components: FedMLAC-Update and FedMLAC-Aggregation. The first component, FedMLAC-Update, takes place on the federated clients and aims to update both the local client AC models and the global Plug-in AC model through a mutual learning mechanism. This process facilitates the transfer of global federated knowledge to the client AC models while simultaneously updating the global Plug-in AC model with new local knowledge.
The second component, FedMLAC-Aggregation, operates on the server side. It aggregates the updated Plug-in AC models uploaded by the clients, resulting in an updated global Plug-in AC model. This updated global AC model serves as an auxiliary tool for the subsequent round of client AC model updates, as illustrated in Fig. 1.
Figure 1: Overview of the proposed FedMLAC framework. The framework consists of four key steps: ① each client downloads the globally shared lightweight Plug-in model from the server; ② local mutual learning is performed between the personalized client model and the Plug-in model to enable bidirectional knowledge transfer; ③ the updated Plug-in models are uploaded to the server; ④ the server aggregates these models using the proposed Layer-wise Pruning Aggregation (LPA) strategy to form a new global Plug-in model. This iterative process enables continuous learning across heterogeneous clients while improving robustness to noisy or non-IID data.
4.2 FedMLAC Update
Algorithm 1 FedMLAC-Update
Input: Server-shared global Plug-in model $\theta_{t}^{k}$ and client model $\omega_{t}^{k}$
Output: Updated $\theta_{t+1}^{k}$ and $\omega_{t+1}^{k}$
1:procedure FedMLAC-Update($\theta_{t}^{k}$,$k$,$t$):
2:     $\triangleright$ Run on the client $k$
3:     for each epoch $i$ from $1$ to $E$ do
4:         for each batch $\mathcal{B}\in\mathcal{D}_{k}$ do
5:              Compute loss $\mathcal{L}^{k}(\mathcal{B};\omega^{k}_{t},\theta_{t}^{k})$ using Eq. (7)
6:              $\triangleright$ Client model training loss
7:              Update client model parameters using SGD:
8:              $\omega^{k}_{t+(i+1)}\leftarrow\omega^{k}_{t+(i)}-\eta\nabla\mathcal{L}^{k}(\mathcal{B};\omega^{k}_{t+(i)},\theta^{k}_{t+(i)})$
9:              $\triangleright$ $\eta$ is the learning rate.
10:              Compute loss $\mathcal{L}^{g}(\mathcal{B};\omega^{k}_{t},\theta^{k}_{t})$ using Eq. (8)
11:              $\triangleright$ Plug-in model training loss
12:              Update Plug-in model parameters using SGD:
13:              $\theta^{k}_{t+(i+1)}\leftarrow\theta^{k}_{t+(i)}-\eta\nabla\mathcal{L}^{g}(\mathcal{B};\omega^{k}_{t+(i)},\theta^{k}_{t+(i)})$
14:         end for
15:     end for
16:     return $\theta^{k}_{t+1}$
17:end procedure
Unlike traditional FedAC frameworks, our FedMLAC framework allows each client to have its own personalized AC model, making the framework more aligned with real-world applications. Additionally, each client receives a global Plug-in AC model from the server, which participates in the training process alongside the local AC model. The specific training objectives for these two AC models are described below.
4.2.1 Local AC Model Update
At the beginning of each federated communication round, the local AC model $\omega^{k}_{t}$ undergoes training and parameter updates on the private dataset. Consequently, its training objective must include an AC task-specific loss function, such as the commonly used CE loss, which is defined as follows:
$\mathcal{L}^{k}_{CE}((x,y);\omega^{k}_{t})=CE(F_{k}(x;\omega^{k}_{t}),y),$ (5)
where $(x,y)\in D_{k}$, and $F_{k}$ represents the $k^{th}$ client AC model.
In addition to the AC task-specific loss, to obtain a more adaptable and robust local AC model, each client participating in FL should acquire knowledge about the AC data distributions of other clients. This can be achieved by treating the global Plug-in AC model $\theta^{k}_{t}$ as a teacher and the local AC model as a student, enabling the local AC model to effectively learn global knowledge. Specifically, a KD loss is added to the local AC model’s training objective, defined as:
$\mathcal{L}^{k}_{KL}(x;\omega^{k}_{t},\theta^{k}_{t})=KL\left(F_{g}(x;\theta^{k}_{t}),F_{k}(x;\omega^{k}_{t}))\right),$ (6)
wherein $KL(\cdot,\cdot)$ denotes the KL divergence loss. By combining Eqs. (5) and (6), we reformulate the client AC model training objective function in the $(t+1)^{th}$ communication round as:
$_{\omega^{k}_{t+1}}^{min}\mathcal{L}^{k}((x,y);\omega^{k}_{t},\theta_{t}^{k})=\alpha\mathcal{L}^{k}_{CE}+(1-\alpha)\mathcal{L}_{KL}^{k},$ (7)
where $\alpha$ is a balancing parameter that controls the relative weight of the two loss terms.
4.2.2 Global Plug-in Model Update
For the global Plug-in model, in addition to serving as a carrier of global knowledge, it also participates in training on each client to learn the local AC data knowledge of that client. This enables the Plug-in model to update its existing knowledge based on local client data. In this case, the local client AC model acts as the teacher, and the Plug-in model serves as the student. Consequently, the training objective for updating the Plug-in model includes a knowledge distillation loss, defined as:
$\mathcal{L}^{g}(x;\omega^{k}_{t},\theta^{k}_{t})=KL\left(F_{k}(x;\omega^{k}_{t}),F_{g}(x;\theta^{k}_{t}))\right).$ (8)
This loss ensures that the Plug-in model effectively integrates client-specific AC data knowledge, enhancing its adaptability and robustness for global FL. The entire FedMLAC-Update process is detailed in Algorithm 1.
4.3 FedMLAC Aggregation
Once the global Plug-in model on each client completes its local knowledge update, it is uploaded to the server for aggregation to generate a new global Plug-in model. In traditional federated aggregation (e.g., Eq. (3)), all client-uploaded Plug-in models $\{\theta_{t+1}^{k}\}_{k\in S_{t}}$ are directly combined using data-size weighted averaging. However, under non-IID data distributions or adversarial conditions (e.g., noisy inputs, label-flipping, or poisoned data), some local Plug-in models may introduce substantial deviations from the true global optimum. These deviations manifest as abnormally large or small parameters in specific layers, thereby biasing the global Plug-in model when such updates are indiscriminately aggregated.
To address this, we introduce a robust layer-wise aggregation strategy named Layer-wise Pruning Aggregation (LPA), which aims to eliminate extreme deviations by pruning abnormal client updates at each individual layer before aggregation. This strategy improves robustness against both statistical heterogeneity and malicious updates by identifying outlier layers that exhibit inconsistent behavior relative to the majority.
Layer-wise Similarity Computation
Let $\theta_{t+1}^{k}=\{\theta_{t+1}^{k,l}\}_{l=1}^{L}$ denote the $k^{th}$ client’s Plug-in AC model at round $t+1$, where $\theta_{t+1}^{k,l}$ is the parameter vector of the $l^{th}$ layer. For each layer $l$, we define the average model across clients as:
$\bar{\theta}_{t+1}^{l}=\frac{1}{|S_{t}|}\sum_{k\in S_{t}}\theta_{t+1}^{k,l}.$ (9)
We then measure the deviation of each Plug-in model’s layer parameters from the average using the $\ell_{2}$ norm:
$\delta^{k,l}=\|\theta_{t+1}^{k,l}-\bar{\theta}_{t+1}^{l}\|_{2}.$ (10)
This value reflects how much Plug-in model $k$’s layer-$l$ parameters deviate from the average behavior of the cohort. A high deviation may indicate either statistical heterogeneity or an adversarial anomaly.
Algorithm 2 FedMLAC-Aggregation
Input: Client-uploaded Plug-in models $\{\theta_{t+1}^{k}\}_{k\in S_{t}}$
Output: Aggregated global Plug-in model $\theta_{t+1}^{g}$
1:procedure FedMLAC-Aggregation($\{\theta_{t+1}^{k}\}_{k\in S_{t}}$)
2:     $\triangleright$ Run on the server
3:     for each model $k\in S_{t}$ do
4:         for each model layer $l$ from $1$ to $L$ do
5:              Compute $\bar{\theta}_{t+1}^{l}$ using Eq. (9)
6:              $\triangleright$ average model layer across $\{\theta_{t+1}^{k}\}_{k\in S_{t}}$)
7:              Compute $\delta^{k,l}$ using Eq. (10)
8:              $\triangleright$ $l^{th}$ model layer deviation
9:              Collect layer deviations: $\bar{\mathcal{T}}_{l}\leftarrow\delta^{k,l}$
10:         end for
11:     end for
12:     Sort and prune $\bar{\mathcal{T}}_{l}$ as $\mathcal{T}_{l}$ according to Eq. (11)
13:     $\triangleright$ prune high-deviation layers
14:     Compute $\theta_{t+1}^{g,l}=\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\theta_{t+1}^{k,l}$
15:     $\triangleright$ $l^{th}$ model layer-wise pruning aggregation
16:     Form the new global Plug-in model $\theta_{t+1}^{g}=\{\theta_{t+1}^{g,l}\}_{l=1}^{L}$
17:end procedure
Pruning High-Deviation Layers
To mitigate the influence of outlier updates, we prune the Plug-in models that exhibit extreme deviations at each layer. Specifically, for each layer $l$, we first compute the deviation scores $\delta^{k,l}$ for all clients $k\in S_{t}$ using Eq. (10). These scores are then sorted in ascending order, and both the top-$\lfloor v_{h}\cdot|S_{t}|\rfloor$ and bottom-$\lfloor v_{l}\cdot|S_{t}|\rfloor$ clients are removed from consideration. The remaining clients form the trusted subset $\mathcal{T}_{l}$ used for layer-$l$ aggregation, formally defined as:
$\mathcal{T}_{l}=\left\{k\in S_{t}\;\middle|\;\delta^{k,l}\in\text{Middle}\left(\left\{\delta^{k,l}\right\}_{k\in S_{t}},v_{l},v_{h}\right)\right\},$ (11)
where the operator $\text{Middle}(\cdot,v_{l},v_{h})$ returns the central subset of deviations after removing the largest $\lfloor v_{h}\cdot|S_{t}|\rfloor$ and smallest $\lfloor v_{l}\cdot|S_{t}|\rfloor$ elements. Accordingly, the size of the trusted set is:
$|\mathcal{T}_{l}|=|S_{t}|-\lfloor v_{h}\cdot|S_{t}|\rfloor-\lfloor v_{l}\cdot|S_{t}|\rfloor.$ (12)
This pruning strategy helps reduce the impact of statistical heterogeneity or adversarial perturbations by excluding model updates with extreme deviations prior to aggregation.
The pruned, layer-wise aggregation is then performed as:
$\theta_{t+1}^{g,l}=\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\theta_{t+1}^{k,l},$ (13)
where $n_{k}$ is the dataset size of client $k$ and $n_{\mathcal{T}_{l}}=\sum_{k\in\mathcal{T}_{l}}n_{k}$ is the total data size of the trusted clients for layer $l$. The final global Plug-in model $\theta_{t+1}^{g}$ is then reconstructed by concatenating all the aggregated layers: $\theta_{t+1}^{g}=\{\theta_{t+1}^{g,l}\}_{l=1}^{L}$. The specific implementation procedure of FedMLAC-Aggregation is provided in Algorithm 2.
4.4 Disscusion
FedMLAC embodies a unified design philosophy that bridges personalization, scalability, and robustness in FedAC. Its core strength lies in decoupling global knowledge aggregation from personalized model learning through a collaborative Plug-in mechanism. This architectural separation not only supports model heterogeneity, but also allows clients to retain autonomy over their model structures and local objectives—crucial in diverse AC scenarios where tasks, devices, and environments differ significantly.
Moreover, by framing mutual learning as bidirectional knowledge distillation, FedMLAC integrates both generalization and adaptation within each training round. This mitigates the classic trade-off between convergence stability and model personalization, which often plagues traditional FL pipelines. The LPA component further ensures resilience to noisy or corrupted updates, particularly relevant for audio data subject to ambiguous labels, device noise, or inconsistent recording quality.
5 Convergence Analysis of FedMLAC
This section presents a rigorous convergence analysis of the FedMLAC framework, focusing on the global Plug-in model’s convergence under the challenges of data heterogeneity, model heterogeneity, and data poisoning. We formalize the theoretical assumptions, derive the convergence properties, and present a detailed proof to demonstrate that FedMLAC achieves convergence to a stationary point with a quantifiable rate.
5.1 Theoretical Assumptions
To facilitate the convergence analysis, we adopt the following standard assumptions, commonly used in FL literature:
Assumption 1 (Smoothness)
The local loss functions $\mathcal{L}^{k}(\omega)$ and $\mathcal{L}^{g}(\theta)$ are $L$-smooth, i.e., for all $\omega,\omega^{\prime},\theta,\theta^{\prime}$:
$\|\nabla\mathcal{L}^{k}(\omega)-\nabla\mathcal{L}^{k}(\omega^{\prime})\|\leq L\|\omega-\omega^{\prime}\|,\quad\|\nabla\mathcal{L}^{g}(\theta)-\nabla\mathcal{L}^{g}(\theta^{\prime})\|\leq L\|\theta-\theta^{\prime}\|.$ (14)
Assumption 2 (Bounded Gradients)
The gradients of the local loss $\mathcal{L}^{k}(\omega)$ and the Plug-in model loss $\mathcal{L}^{g}(\theta)$ are bounded:
$\|\nabla\mathcal{L}^{k}(\omega)\|^{2}\leq G^{2},\quad\|\nabla\mathcal{L}^{g}(\theta)\|^{2}\leq G^{2}.$ (15)
Assumption 3 (Bounded Data Heterogeneity)
The local loss gradients exhibit bounded heterogeneity with respect to the global loss:
$\mathbb{E}[\|\nabla\mathcal{L}^{k}(\theta)-\nabla\mathcal{L}(\theta)\|^{2}]\leq\sigma^{2},$ (16)
where $\sigma^{2}$ quantifies the degree of non-IID data across clients.
Assumption 4 (Bounded Stochastic Gradient Variance)
The stochastic gradients computed on mini-batches $\mathcal{B}\subset\mathcal{D}_{k}$ are unbiased with bounded variance:
$\mathbb{E}[\nabla\mathcal{L}^{k}(\omega;\mathcal{B})]=\nabla\mathcal{L}^{k}(\omega),\quad\mathbb{E}[\|\nabla\mathcal{L}^{k}(\omega;\mathcal{B})-\nabla\mathcal{L}^{k}(\omega)\|^{2}]\leq\sigma_{g}^{2},$ (17)
$\mathbb{E}[\nabla\mathcal{L}^{g}(\theta;\mathcal{B})]=\nabla\mathcal{L}^{g}(\theta),\quad\mathbb{E}[\|\nabla\mathcal{L}^{g}(\theta;\mathcal{B})-\nabla\mathcal{L}^{g}(\theta)\|^{2}]\leq\sigma_{g}^{2}.$ (18)
Assumption 5 (LPA Effectiveness)
The LPA strategy ensures that the parameters of the trusted subset $\mathcal{T}_{l}$ for each layer $l$ have bounded deviation from the average:
$\mathbb{E}[\|\theta_{t+1}^{k,l}-\bar{\theta}_{t+1}^{l}\|^{2}]\leq\delta^{2},\quad k\in\mathcal{T}_{l},$ (19)
where $\bar{\theta}_{t+1}^{l}=\frac{1}{|S_{t}|}\sum_{k\in S_{t}}\theta_{t+1}^{k,l}$, and $\delta^{2}$ is the upper bound on the deviation after pruning.
Assumption 6 (Bounded Loss)
The global loss $\mathcal{L}(\theta)$ is bounded below:
$\mathcal{L}(\theta)\geq\mathcal{L}^{*}>-\infty.$ (20)
These assumptions ensure that the loss functions are well-behaved, gradients are manageable, and the LPA strategy effectively mitigates outliers due to data poisoning or heterogeneity. While we assume non-convexity (common for deep learning models), the smoothness and boundedness assumptions provide sufficient structure for convergence analysis.
5.2 Convergence Analysis
We analyze the convergence of the global Plug-in model $\theta^{g}$ to a stationary point of the global objective $\mathcal{L}(\theta^{g})$. The analysis proceeds in three steps: (1) bounding the loss decrease during local updates of $\omega^{k}$ and $\theta^{k}$, (2) analyzing the effect of LPA on global aggregation, and (3) deriving the overall convergence rate.
5.2.1 Local Update Analysis
For each client $k$, the local model $\omega^{k}$ is updated via SGD as below:
$\omega_{t+(i+1)}^{k}=\omega_{t+(i)}^{k}-\eta\nabla\mathcal{L}^{k}(\mathcal{B};\omega_{t+(i)}^{k},\theta_{t+(i)}^{k}),$ (21)
where $\eta$ is the learning rate, and $i=0,1,\ldots,E-1$ indexes local epochs. Since $\mathcal{L}^{k}$ is $L$-smooth (Assumption 1), we apply the descent lemma:
$\mathbb{E}[\mathcal{L}^{k}(\omega_{t+(i+1)}^{k})]$ $\leq\mathcal{L}^{k}(\omega_{t+(i)}^{k})-\eta\left\langle\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k}),\mathbb{E}\left[\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k};\mathcal{B})\right]\right\rangle$
$\quad+\frac{L\eta^{2}}{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k};\mathcal{B})\right\|^{2}\right].$ (22)
By Assumption 4, the stochastic gradient is unbiased: $\mathbb{E}[\nabla\mathcal{L}^{k}(\omega;\mathcal{B})]=\nabla\mathcal{L}^{k}(\omega)$. The gradient variance bound gives:
$\mathbb{E}[\|\nabla\mathcal{L}^{k}(\omega;\mathcal{B})\|^{2}]$ $=\mathbb{E}[\|\nabla\mathcal{L}^{k}(\omega;\mathcal{B})-\nabla\mathcal{L}^{k}(\omega)+\nabla\mathcal{L}^{k}(\omega)\|^{2}]$
$\leq\|\nabla\mathcal{L}^{k}(\omega)\|^{2}+\sigma_{g}^{2}$
$\leq G^{2}+\sigma_{g}^{2}.$ (23)
where the last inequality follows from Assumption 2. Thus:
$\mathbb{E}[\mathcal{L}^{k}(\omega_{t+(i+1)}^{k})]\leq\mathcal{L}^{k}(\omega_{t+(i)}^{k})-\eta\|\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}}{2}(G^{2}+\sigma_{g}^{2}).$ (24)
Summing over $E$ local epochs:
$\mathbb{E}[\mathcal{L}^{k}(\omega_{t+1}^{k})]\leq\mathcal{L}^{k}(\omega_{t}^{k})-\eta\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}E}{2}(G^{2}+\sigma_{g}^{2}).$ (25)
Similarly, for the Plug-in model $\theta^{k}$, updated via:
$\theta_{t+(i+1)}^{k}=\theta_{t+(i)}^{k}-\eta\nabla\mathcal{L}^{g}(\mathcal{B};\omega_{t+(i)}^{k},\theta_{t+(i)}^{k}),$ (26)
we obtain:
$\mathbb{E}[\mathcal{L}^{g}(\theta_{t+(i+1)}^{k})]\leq\mathcal{L}^{g}(\theta_{t+(i)}^{k})-\eta\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}}{2}(G^{2}+\sigma_{g}^{2}).$ (27)
Summing over $E$ epochs:
$\mathbb{E}[\mathcal{L}^{g}(\theta_{t+1}^{k})]\leq\mathcal{L}^{g}(\theta_{t}^{k})-\eta\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}E}{2}(G^{2}+\sigma_{g}^{2}).$ (28)
The mutual learning mechanism, through the KL divergence terms in Eqs. (8) and  (7), aligns the local model $\omega^{k}$ with the global knowledge in $\theta^{k}$ and vice versa, reducing the impact of data heterogeneity by regularizing the local optimization direction.
5.2.2 Global Aggregation with LPA
In the FedMLAC-Aggregation step, the server aggregates the client-uploaded Plug-in models $\{\theta_{t+1}^{k}\}_{k\in S_{t}}$ using LPA. For each layer $l$, LPA computes the average parameters:
$\bar{\theta}_{t+1}^{l}=\frac{1}{|S_{t}|}\sum_{k\in S_{t}}\theta_{t+1}^{k,l},$ (29)
and the deviation for each client $k$:
$\delta^{k,l}=\|\theta_{t+1}^{k,l}-\bar{\theta}_{t+1}^{l}\|_{2}.$ (30)
The trusted subset $\mathcal{T}_{l}$ excludes the top $\lfloor v_{h}\cdot|S_{t}|\rfloor$ and bottom $\lfloor v_{l}\cdot|S_{t}|\rfloor$ clients based on $\delta^{k,l}$, and the global layer parameters are computed as:
$\theta_{t+1}^{g,l}=\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\theta_{t+1}^{k,l},$ (31)
where $n_{\mathcal{T}_{l}}=\sum_{k\in\mathcal{T}_{l}}n_{k}$. The global Plug-in model is $\theta_{t+1}^{g}=\{\theta_{t+1}^{g,l}\}_{l=1}^{L}$.
To analyze the effect of LPA, we consider the deviation of the aggregated model from the true global optimum. By Assumption 5, the parameters in $\mathcal{T}_{l}$ satisfy $\mathbb{E}[\|\theta_{t+1}^{k,l}-\bar{\theta}_{t+1}^{l}\|^{2}]\leq\delta^{2}$. The global loss after aggregation can be bounded using the smoothness of $\mathcal{L}$ (Assumption 1):
$\mathbb{E}[\mathcal{L}(\theta_{t+1}^{g})]\leq\mathcal{L}(\theta_{t}^{g})+\langle\nabla\mathcal{L}(\theta_{t}^{g}),\mathbb{E}[\theta_{t+1}^{g}-\theta_{t}^{g}]\rangle+\frac{L}{2}\mathbb{E}[\|\theta_{t+1}^{g}-\theta_{t}^{g}\|^{2}].$ (32)
The expected update direction is:
$\mathbb{E}[\theta_{t+1}^{g}-\theta_{t}^{g}]=\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}(\theta_{t+1}^{k,l}-\theta_{t}^{g,l}).$ (33)
Since $\theta_{t+1}^{k,l}$ results from $E$ local SGD steps, we approximate:
$\theta_{t+1}^{k,l}-\theta_{t}^{g,l}\approx-\eta\sum_{i=0}^{E-1}\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k};\mathcal{B}).$ (34)
Thus:
$\mathbb{E}[\theta_{t+1}^{g}-\theta_{t}^{g}]\approx-\eta\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k}).$ (35)
The second term, $\mathbb{E}[\|\theta_{t+1}^{g}-\theta_{t}^{g}\|^{2}]$, is influenced by the variance of local updates and the pruning effect of LPA:
$\mathbb{E}[\|\theta_{t+1}^{g}-\theta_{t}^{g}\|^{2}]\leq\eta^{2}E(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}),$ (36)
where $\sigma^{2}$ accounts for data heterogeneity (Assumption 3), and $\delta^{2}$ accounts for residual deviations after LPA (Assumption 5). Combining these, the global loss decrease is:
$\mathbb{E}[\mathcal{L}(\theta_{t+1}^{g})]\leq\mathcal{L}(\theta_{t}^{g})-\eta\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}E}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (37)
5.2.3 Convergence Theorem
We now present the main convergence result for FedMLAC.
Theorem 1
Under Assumptions 16, with learning rate $\eta<\frac{1}{L}$, after $T$ communication rounds, the expected gradient norm of the global Plug-in model satisfies:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t}^{g})\|^{2}]\leq\frac{\mathcal{L}(\theta_{0}^{g})-\mathcal{L}^{*}}{\eta TE}+\frac{L\eta E}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (38)
Summing the global loss decrease (37) over $T$ rounds:
$\sum_{t=0}^{T-1}\mathbb{E}[\mathcal{L}(\theta_{t+1}^{g})-\mathcal{L}(\theta_{t}^{g})]$ $\leq-\eta\sum_{t=0}^{T-1}\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\left\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\right\|^{2}$
$\quad+\frac{L\eta^{2}ET}{2}\left(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}\right).$ (39)
By Assumption 6, the left-hand side is:
$\sum_{t=0}^{T-1}(\mathcal{L}(\theta_{t+1}^{g})-\mathcal{L}(\theta_{t}^{g}))=\mathcal{L}(\theta_{T}^{g})-\mathcal{L}(\theta_{0}^{g})\geq\mathcal{L}^{*}-\mathcal{L}(\theta_{0}^{g}).$ (40)
Rearranging:
$\eta\sum_{t=0}^{T-1}\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}\leq\mathcal{L}(\theta_{0}^{g})-\mathcal{L}^{*}+\frac{L\eta^{2}ET}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (41)
Assuming $\theta_{t+(i)}^{k}\approx\theta_{t}^{g}$ for small $\eta$, we approximate:
$\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}\approx E\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t}^{g})\|^{2}].$ (42)
Dividing by $\eta TE$, we obtain:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t}^{g})\|^{2}]\leq\frac{\mathcal{L}(\theta_{0}^{g})-\mathcal{L}^{*}}{\eta TE}+\frac{L\eta E}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (43)
5.3 Discussion
The convergence rate of FedMLAC is $\mathcal{O}(1/\sqrt{T})$ when $\eta\propto 1/\sqrt{T}$, which is standard for non-convex optimization in FL. The mutual learning mechanism aligns local and global objectives, reducing the impact of $\sigma^{2}$ by regularizing local updates with global knowledge. The LPA strategy ensures robustness against data poisoning by bounding $\delta^{2}$, as it prunes anomalous updates that could destabilize aggregation. The lightweight Plug-in model reduces communication overhead, enhancing scalability in heterogeneous environments.
6 Experiments
In this section, we conduct comprehensive experiments to evaluate the effectiveness of the proposed FedMLAC framework. Our evaluation spans four benchmark datasets covering both speech and non-speech audio classification tasks. We assess FedMLAC in terms of classification accuracy, robustness to noisy and heterogeneous data, and personalization capability, comparing it against three representative state-of-the-art FL baselines.
6.1 Experimental Settings
Table 2: Summary of federated settings across benchmark datasets, including audio classification tasks, partition strategies, and data statistics.
Dataset AC Task Partition Strategy #Clients #Samples #Classes
GSC Keyword Spotting Speaker ID 2,618 105,829 35
IEMOCAP Emotion Recognition Actor ID 10 2,943 4
CREMA-D Emotion Recognition Speaker ID 91 4,798 4
Urban Sound Sound Event Detection Dirichlet Distribution 50 8,732 10
Datasets. We evaluate FedMLAC on four publicly available datasets widely adopted in federated audio classification (FedAC), as summarized in Table 2. (1) Google Speech Commands (GSC) [68] is a large-scale benchmark for keyword spotting, comprising over 100,000 short utterances covering 35 command words. (2) Urban Sound [69] includes environmental audio samples spanning 10 everyday sound event classes, such as sirens, drilling, and dog barking. (3) IEMOCAP [70] is a multimodal dataset for emotion recognition. In this work, we focus on its audio-only component, selecting four commonly used emotions from the improvised sessions. (4) CREMA-D [71] features speech clips performed by actors under emotional prompts. We consider the audio channel and select four target emotions consistent with prior studies. To ensure fair comparison across datasets, we apply consistent preprocessing strategies including silence trimming, segmentation, and acoustic feature extraction (e.g., Mel spectrograms or pretrained APC embeddings), following the protocol proposed in [72].
Data Partition. To simulate realistic client distributions in federated settings, we adopt dataset-specific non-IID partitioning strategies. For the Google Speech Commands and CREMA-D datasets, we assign samples to clients based on speaker identity, ensuring that each client contains utterances from a single speaker. This setup reflects naturally occurring data silos in speech-based systems. In the case of IEMOCAP, we divide data by actor identity, with each client corresponding to the recordings of one actor. This preserves speaker consistency within clients while exposing variation across clients. For the Urban Sound dataset, we follow a label distribution skewing strategy using the Dirichlet allocation method, as commonly adopted in FL studies [73]. Specifically, we vary the Dirichlet concentration parameter $\alpha\in\{0.1,0.3,0.5,1.0\}$ to control the degree of data heterogeneity. Smaller values of $\alpha$ lead to greater data heterogeneity across clients, whereas larger values result in more balanced and uniform data distributions.
AC Models. To support AC in both homogeneous and heterogeneous FL settings, we construct a suite of Convolutional Recurrent Neural Network (CRNN) architectures with varying levels of complexity as listed in Table 3. All models share a consistent input-output format and are compatible with the same audio preprocessing pipeline (e.g., Mel-spectrogram or APC features), ensuring fair performance comparison across settings.
In the homogeneous setting, all clients employ the same model architecture, namely CRNN-Base, which consists of two 1D convolutional layers (64 filters) followed by ReLU activation, max-pooling, and dropout for feature extraction, and a bidirectional GRU with 128 hidden units for temporal modeling. The extracted features are fed into a fully connected output layer for final prediction. To enable mutual knowledge transfer, each client is also equipped with a lightweight Plug-in model, instantiated as CRNN-Lite, which adopts a compact architecture with two 1D convolutional layers (32 filters) and a unidirectional GRU with 64 hidden units.
In the heterogeneous setting, to emulate diverse client capabilities, each client is randomly assigned one of the five CRNN variants listed in Table 3. These models differ in convolutional depth and GRU configuration but maintain similar classification capacity, enabling controlled architectural heterogeneity across federated clients.
Table 3: CRNN model variants used in homogeneous and heterogeneous client settings.
Model Name Conv Layers RNN Type Hidden Units
CRNN-Tiny 1 × Conv1D (16 filters) GRU 32
CRNN-Lite 2 × Conv1D (32 filters) GRU 64
CRNN-Mid 3 × Conv1D (32 filters) GRU 64
CRNN-Base 2 × Conv1D (64 filters) Bi-GRU 128
CRNN-Deep 3 × Conv1D (64→128 filters) Bi-GRU 128
Hyperparameters. In our FedAC experiments, we adopt the following training configurations to balance convergence speed and communication efficiency. The maximum number of communication rounds is set to 5000. Each client performs one local epoch per round with a batch size of 16. The learning rate is fixed at 0.01. Unless otherwise specified, 20% of clients are randomly selected to participate in each communication round.
Metric. We evaluate performance on the Google Speech Commands dataset using classification accuracy, averaged over five independent runs with different random seeds. For the other three datasets, we adopt the F1 score as the evaluation metric, reporting the mean and standard deviation across all runs. F1 score is used to account for class imbalance, which is common in emotion recognition and sound event classification tasks.
Baselines. We compare our method with three widely used FL algorithms: FedAvg [10], FedProx [61], and FedOPT [62]. These methods serve as strong general-purpose baselines and are commonly adopted across diverse FL tasks, including audio applications. We do not include FedAC-specific approaches due to the lack of standardized implementations and their reliance on incompatible datasets or architectures, which hinders fair comparison.
6.2 Performance Under Homogeneous Models
To assess the effectiveness of FedMLAC under homogeneous model conditions, we evaluate its performance against three representative FL baselines: FedAvg, FedProx, and FedOPT. In each communication round, only 20% of clients are randomly selected to participate, except for the IEMOCAP dataset, where all clients are activated due to its small client population. Unlike FedMLAC, which maintains personalized local models and exchanges lightweight Plug-in models with the server, baseline methods directly share full client models for aggregation. All methods are evaluated under identical training protocols, and for fair comparison, we report the averaged performance of all local client models at the end of training. This ensures that the comparison reflects the collective quality of personalized models across the federation, rather than relying solely on a global proxy model.
Table 4: Performance comparison of different FL methods (Active Ratio = 20% except IEMOCAP) under homogeneous models.
Dataset
#Clients
(Active Ratio)
FedAvg FedProx FedOPT FedMLAC
GSC 524 (20%) 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 84.59$\pm$0.19
IEMOCAP 10 (100%) 48.54$\pm$6.59 48.89$\pm$6.39 50.01$\pm$6.77 51.97$\pm$5.81
CREMA-D 18 (20%) 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 63.04$\pm$3.42
Urban Sound
$\alpha=0.1$
10 (20%) 50.43$\pm$4.73 50.62$\pm$5.58 48.92$\pm$5.83 52.86$\pm$4.06
Urban Sound
$\alpha=0.3$
10 (20%) 51.83$\pm$4.42 52.07$\pm$4.34 50.36$\pm$5.61 54.29$\pm$4.17
Urban Sound
$\alpha=0.5$
10 (20%) 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 55.61$\pm$3.83
Urban Sound
$\alpha=1.0$
10 (20%) 54.64$\pm$4.40 54.89$\pm$3.54 53.03$\pm$4.93 56.95$\pm$3.65
As shown in Table 4, FedMLAC consistently outperforms all baseline methods across datasets and varying degrees of data heterogeneity. On Google Speech Commands, a large-scale and relatively balanced dataset, FedMLAC surpasses FedOPT by 1.28%, demonstrating its ability to effectively integrate global knowledge even under limited client participation. For IEMOCAP, which poses challenges due to high inter-speaker variability and data scarcity, FedMLAC yields up to a 3.43% improvement over FedAvg and 3.08% over FedProx, highlighting its robustness in low-resource, high-variance environments.
On the CREMA-D dataset, despite reduced client activation, FedMLAC achieves superior performance, reflecting its resilience to partial participation and speaker-specific skew. In the Urban Sound experiments, where the Dirichlet parameter $\alpha$ controls the degree of heterogeneity of federated data distributions, all methods show improved accuracy as $\alpha$ increases. However, FedMLAC consistently maintains a clear advantage across all heterogeneity levels, with accuracy rising from 52.86% at $\alpha=0.1$ to 56.95% at $\alpha=1.0$. These results collectively underscore the effectiveness of FedMLAC’s mutual learning mechanism and layer-wise robust aggregation in handling both personalization and non-IID challenges in FedAC.
6.3 Performance Under Heterogeneous Models
In this subsection, we assess the performance of FedMLAC where clients employ heterogeneous local model architectures, a realistic setting in edge computing and cross-device FL. Since conventional FL baselines such as FedAvg, FedProx, and FedOPT require model consistency across clients, they are excluded from this comparison. Instead, we evaluate FedMLAC against its ablated variant (FedMLAC w/o LPA), where the proposed LPA is disabled. All remaining experimental conditions, including client participation ratio and training protocol, are consistent with those in Section 6.2 to ensure fair comparison.
Figure 2: Performance comparison between FedMLAC and FedMLAC without LPA under heterogeneous model settings across four AC datasets. Each bar represents the average accuracy across all clients, with error bars indicating standard deviation. “GSC” denotes Google Speech Commands, and “US” refers to Urban Sound.
As shown in Figure 2, FedMLAC consistently outperforms its variant without LPA across all tested AC datasets and non-IID configurations. On Google Speech Commands, FedMLAC improves accuracy by 1.75%, indicating stable global knowledge transfer even under architectural diversity. In low-resource emotion datasets like IEMOCAP and CREMA-D, where speaker-induced variation is prominent, FedMLAC yields 2.04% and 5.23% accuracy improvements, respectively, suggesting superior robustness against speaker and model shifts. Moreover, on Urban Sound, where Dirichlet-based non-IID data distribution is introduced, FedMLAC achieves steady gains across all levels of heterogeneity. Notably, under the most skewed setting ($\alpha=0.1$), it outperforms its ablated version by 3.82%, demonstrating that LPA plays a key role in filtering noisy updates and stabilizing aggregation under model and data heterogeneity. These results highlight the essential contribution of LPA in achieving generalizable and resilient performance in heterogeneous FedAC systems.
6.4 Performance Under Varying Active Ratio
To evaluate the impact of client participation rates on model performance, we study the behavior of different FL algorithms under varying active client ratios, ranging from 10% to 100%. This setting reflects real-world deployment scenarios where only a subset of clients may participate in each communication round due to resource or connectivity constraints. As client participation directly affects the data diversity seen by the server per round, it plays a critical role in model convergence and generalization. In this experiment, we exclude IEMOCAP due to its limited client pool. All other configurations, including the homogeneous model architecture and training hyperparameters, follow the setup described in Section 6.2.
Table 5: Impact of varying active client ratios (10%, 20%, 50%, and 100%) on the performance of different FL methods under homogeneous models.
Dataset
#Clients
(Active Ratio)
FedAvg FedProx FedOPT FedMLAC
GSC 262 (10%) 78.14$\pm$0.52 78.30$\pm$0.49 81.92$\pm$0.22 82.81$\pm$0.17
524 (20%) 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 84.59$\pm$0.19
1310 (50%) 81.73$\pm$0.58 81.95$\pm$0.35 83.82$\pm$0.31 85.10$\pm$0.21
2618 (100%) 82.58$\pm$0.46 82.73$\pm$0.42 84.84$\pm$0.19 85.76$\pm$0.13
CREMA-D 9 (10%) 59.87$\pm$4.02 60.10$\pm$3.78 58.02$\pm$4.23 61.78$\pm$3.62
18 (20%) 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 63.04$\pm$3.42
45 (50%) 61.72$\pm$3.75 61.85$\pm$4.16 59.52$\pm$4.05 63.93$\pm$2.24
91 (100%) 62.57$\pm$3.16 62.45$\pm$3.50 60.20$\pm$3.90 64.79$\pm$2.08
Urban Sound 5 (10%) 50.65$\pm$5.21 51.18$\pm$5.23 49.12$\pm$5.89 52.74$\pm$4.26
10 (20%) 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 55.61$\pm$3.83
25 (50%) 53.52$\pm$5.02 54.29$\pm$3.83 52.35$\pm$5.26 57.15$\pm$5.24
50 (100%) 55.21$\pm$3.96 55.47$\pm$4.74 53.12$\pm$4.41 58.30$\pm$3.59
As shown in Table 5, increasing the active client ratio generally improves the accuracy of all FL methods, as more local updates contribute to global model optimization. However, FedMLAC consistently achieves the highest performance across all datasets and participation levels, and demonstrates stronger scalability with respect to client inclusion. For instance, on Google Speech Commands, FedMLAC reaches 85.76% accuracy at 100% participation, outperforming FedOPT and FedAvg by 0.92% and 3.18% respectively. On CREMA-D, FedMLAC improves from 61.78% at 10% participation to 64.79% at full participation, maintaining a consistent lead of 2–3% over all baselines. Urban Sound shows a similar trend, where FedMLAC achieves 58.30% at full participation, surpassing other methods by a margin of 2–5%. These results confirm the robustness of FedMLAC in aggregating diverse local knowledge effectively, especially under partial participation, due to its mutual learning mechanism and layer-wise aggregation that adapt to both statistical and structural heterogeneity.
6.5 Ablation Study
To further investigate the individual contributions of each core component in the proposed FedMLAC framework, we conduct a detailed ablation study under homogeneous model settings. Specifically, we evaluate three variants: the complete FedMLAC, a version without mutual learning (w/o ML) that excludes bidirectional knowledge transfer between the local and Plug-in models, and a version without LPA (w/o LPA) that replaces the robust Layer-wise Pruning Aggregation with standard averaging. This analysis aims to disentangle the effect of global knowledge sharing via mutual learning and robustness enhancement via LPA, thereby validating the design motivation behind FedMLAC. All experimental configurations, including model architectures, dataset partitions, and participation rates, are consistent with those described in Section 6.2.
Figure 3: Ablation comparison between FedMLAC and its variant without LPA and ML under heterogeneous model settings across four AC datasets. Each bar represents the average client-level accuracy, and error bars denote standard deviation across clients. “GSC” corresponds to Google Speech Commands, and “US” represents the Urban Sound dataset.
As shown in Figure 3, the full FedMLAC framework achieves the highest accuracy across all evaluated datasets, highlighting its strength in jointly addressing data and model heterogeneity. On Google Speech Commands, removing mutual learning leads to a noticeable 2.23% performance drop, indicating that the absence of global knowledge transfer hampers generalization. Similarly, disabling LPA results in lower accuracy compared to the complete version, especially on CREMA-D and IEMOCAP, where data noise and variability are pronounced. On Urban Sound, the full FedMLAC consistently outperforms both variants. The performance gap widens as heterogeneity increases (lower $\alpha$), demonstrating that LPA effectively mitigates biased local updates. These results confirm that both the mutual learning mechanism and the LPA module are essential for maintaining stability and boosting performance in FedMLAC
6.6 Performance on Noisy Audio Data
This section investigates the robustness of different FL algorithms under noisy audio conditions. To simulate real-world environments with varying levels of background interference, we add white Gaussian noise to the input audio with different signal-to-noise ratio (SNR) levels. An SNR of 100 dB corresponds to the clean, noise-free condition, while 30 dB, 20 dB, and 10 dB represent increasingly severe noise levels. The Urban Sound dataset is evaluated under a moderately heterogeneous setting with $\alpha=0.5$, which introduces realistic client-level variability alongside noise. Table 6 presents the performance variations of four FL methods across all SNR conditions, with values in parentheses indicating the absolute accuracy drop from the clean condition.
Table 6: Performance comparison of different FL methods under varying SNR levels, including the clean condition (100 dB) and decreasing levels of noise (30 dB, 20 dB, 10 dB). Values in parentheses denote absolute accuracy drop from 100 dB. Results on Urban Sound are reported under $\alpha=0.5$.
Dataset SNR FedAvg FedProx FedOPT FedMLAC
GSC 100dB 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 84.59$\pm$0.19
30dB 76.62$\pm$0.61 ($\downarrow$3.57) 74.70$\pm$0.63 ($\downarrow$5.73) 78.85$\pm$0.31 ($\downarrow$4.46) 81.75$\pm$0.16 ($\downarrow$2.84)
20dB 71.24$\pm$0.58 ($\downarrow$8.95) 68.39$\pm$0.55 ($\downarrow$12.04) 74.44$\pm$0.34 ($\downarrow$8.87) 77.28$\pm$0.29 ($\downarrow$7.31)
10dB 61.26$\pm$0.51 ($\downarrow$18.93) 56.84$\pm$0.46 ($\downarrow$23.59) 67.01$\pm$0.27 ($\downarrow$16.30) 71.27$\pm$0.33 ($\downarrow$13.32)
IEMOCAP 100dB 48.54$\pm$6.59 48.89$\pm$6.39 50.01$\pm$6.77 51.97$\pm$5.81
30dB 46.06$\pm$6.52 ($\downarrow$2.48) 43.89$\pm$7.13 ($\downarrow$5.00) 47.48$\pm$6.69 ($\downarrow$2.53) 50.19$\pm$5.91 ($\downarrow$1.78)
20dB 43.39$\pm$6.64 ($\downarrow$5.15) 40.44$\pm$6.47 ($\downarrow$8.45) 44.98$\pm$6.71 ($\downarrow$5.03) 48.01$\pm$4.84 ($\downarrow$3.96)
10dB 40.33$\pm$6.75 ($\downarrow$8.21) 37.10$\pm$6.29 ($\downarrow$11.79) 41.97$\pm$5.55 ($\downarrow$8.04) 45.82$\pm$6.13 ($\downarrow$6.15)
CREMA-D 100dB 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 63.04$\pm$3.42
30dB 58.37$\pm$3.81 ($\downarrow$2.47) 56.01$\pm$3.83 ($\downarrow$5.01) 55.31$\pm$5.11 ($\downarrow$3.16) 61.55$\pm$3.61 ($\downarrow$1.49)
20dB 54.23$\pm$4.07 ($\downarrow$6.61) 51.91$\pm$3.96 ($\downarrow$9.11) 50.53$\pm$4.52 ($\downarrow$7.94) 58.02$\pm$3.05 ($\downarrow$5.02)
10dB 50.01$\pm$3.62 ($\downarrow$10.83) 46.23$\pm$4.08 ($\downarrow$14.79) 47.13$\pm$4.60 ($\downarrow$11.34) 54.21$\pm$3.84 ($\downarrow$8.83)
Urban Sound 100dB 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 55.61$\pm$3.83
30dB 45.21$\pm$4.34 ($\downarrow$8.22) 43.22$\pm$4.16 ($\downarrow$10.39) 45.31$\pm$6.10 ($\downarrow$6.46) 51.23$\pm$3.96 ($\downarrow$4.38)
20dB 40.02$\pm$4.46 ($\downarrow$13.41) 37.01$\pm$4.32 ($\downarrow$16.60) 40.23$\pm$5.56 ($\downarrow$11.54) 47.32$\pm$3.70 ($\downarrow$8.29)
10dB 32.01$\pm$3.66 ($\downarrow$21.42) 28.83$\pm$4.13 ($\downarrow$24.78) 34.51$\pm$5.69 ($\downarrow$17.26) 43.03$\pm$4.13 ($\downarrow$12.58)
Across all datasets, FedMLAC consistently outperforms baseline methods, achieving both strong accuracy in clean conditions and superior robustness under varying noise levels. For example, on the GSC dataset, FedMLAC yields 81.75% at 30 dB with only a 2.84% drop from the clean setting (100 dB: 84.59%), compared to larger degradations in FedAvg (3.57%), FedProx (5.73%), and FedOPT (4.46%). This trend continues across IEMOCAP and CREMA-D, where FedMLAC exhibits the smallest performance degradation at all SNR levels. At 30 dB, FedMLAC experiences only a 1.78% drop on IEMOCAP and 1.49% on CREMA-D, while FedProx degrades by 5.00% and 5.01%, respectively. Even under severe noise (10 dB), FedMLAC remains more stable, e.g., dropping just 6.15% on IEMOCAP and 8.83% on CREMA-D, significantly less than other methods.
The Urban Sound dataset presents the most pronounced challenge due to its environmental variability. While all methods degrade substantially at low SNR, FedMLAC shows remarkable resilience. At 10 dB, it drops only 12.58%, compared to 21.42% for FedAvg, 24.78% for FedProx, and 17.26% for FedOPT. These results collectively underscore the practical value of FedMLAC, not only for delivering high accuracy in clean conditions but also for preserving performance in real-world noisy federated environments where audio data heterogeneity and corruption are prevalent.
6.7 Performance on Audio Data with Label Errors
This subsection aims to examine how different FL methods perform under varying levels of label noise in AC scenarios. To emulate realistic annotation imperfections, we introduce varying error rates (E.R. = 0.1, 0.3, 0.5) into the training data while keeping the test sets noise-free. Experimental settings remain consistent with earlier sections, including partial client participation (20%) for most datasets and full participation for IEMOCAP. As shown in Table 7, we compare the performance degradation patterns of all methods under noisy supervision, allowing us to assess their robustness across diverse acoustic tasks.
Table 7: Performance comparison of different FL methods under varying label error rates (E.R.), including the clean condition (E.R=0.0) and increasing levels of E.R. (0.1, 0.3, 0.5). Values in parentheses denote absolute performance drop from E.R=0.0. Results on Urban Sound are reported under $\alpha=0.5$.
Dataset E.R. FedAvg FedProx FedOPT FedMLAC
GSC 0.0 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 84.59$\pm$0.19
0.1 79.29$\pm$0.42 ($\downarrow$0.90) 77.59$\pm$0.65 ($\downarrow$2.84) 82.67$\pm$0.75 ($\downarrow$0.64) 84.07$\pm$0.68 ($\downarrow$0.52)
0.3 74.68$\pm$0.66 ($\downarrow$5.51) 76.12$\pm$0.41 ($\downarrow$4.31) 78.19$\pm$0.53 ($\downarrow$5.12) 81.61$\pm$0.25 ($\downarrow$2.98)
0.5 69.70$\pm$0.70 ($\downarrow$10.49) 68.72$\pm$0.64 ($\downarrow$11.71) 74.03$\pm$0.28 ($\downarrow$9.28) 77.06$\pm$0.55 ($\downarrow$7.53)
IEMOCAP 0.0 48.54$\pm$6.59 48.89$\pm$6.39 50.01$\pm$6.77 51.97$\pm$5.81
0.1 47.28$\pm$0.38 ($\downarrow$1.26) 46.18$\pm$0.26 ($\downarrow$2.71) 46.36$\pm$0.66 ($\downarrow$3.65) 51.13$\pm$0.37 ($\downarrow$0.84)
0.3 43.23$\pm$0.39 ($\downarrow$5.31) 43.36$\pm$0.26 ($\downarrow$5.53) 40.14$\pm$0.45 ($\downarrow$9.87) 48.47$\pm$0.45 ($\downarrow$3.23)
0.5 24.24$\pm$0.65 ($\downarrow$24.30) 27.47$\pm$0.38 ($\downarrow$21.42) 30.19$\pm$0.58 ($\downarrow$19.82) 36.28$\pm$0.44 ($\downarrow$15.69)
CREMA-D 0.0 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 63.04$\pm$3.42
0.1 58.23$\pm$0.59 ($\downarrow$2.61) 59.20$\pm$0.64 ($\downarrow$1.82) 57.75$\pm$0.55 ($\downarrow$0.72) 62.06$\pm$0.49 ($\downarrow$0.98)
0.3 55.54$\pm$0.61 ($\downarrow$5.30) 56.34$\pm$0.48 ($\downarrow$4.68) 53.77$\pm$0.27 ($\downarrow$4.70) 60.58$\pm$0.28 ($\downarrow$2.46)
0.5 47.03$\pm$0.31 ($\downarrow$13.81) 46.40$\pm$0.64 ($\downarrow$14.62) 48.24$\pm$0.39 ($\downarrow$10.23) 55.52$\pm$0.69 ($\downarrow$7.52)
Urban Sound 0.0 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 55.61$\pm$3.83
0.1 50.93$\pm$0.26 ($\downarrow$2.50) 52.18$\pm$0.29 ($\downarrow$1.43) 50.51$\pm$0.25 ($\downarrow$1.26) 54.59$\pm$0.71 ($\downarrow$1.02)
0.3 42.66$\pm$0.31 ($\downarrow$10.77) 47.48$\pm$0.30 ($\downarrow$6.13) 40.19$\pm$0.63 ($\downarrow$11.58) 50.36$\pm$0.56 ($\downarrow$5.25)
0.5 35.24$\pm$0.63 ($\downarrow$18.19) 36.74$\pm$0.26 ($\downarrow$16.87) 37.79$\pm$0.51 ($\downarrow$13.98) 44.98$\pm$0.54 ($\downarrow$10.63)
As expected, increasing label noise consistently degrades model performance across all methods and datasets. However, the degree of degradation varies significantly by algorithm and dataset. Notably, FedMLAC demonstrates the highest resilience to label errors in every case. For instance, on the GSC dataset, while FedAvg suffers a 10.49% drop in accuracy at E.R.=0.5, FedMLAC limits this drop to only 7.53%. On IEMOCAP, where the noise sensitivity is more pronounced, FedAvg drops by over 24% and FedProx by 21%, while FedMLAC still retains a notable 36.28% accuracy—reflecting a 15.69% drop, the smallest among all methods.
Similar robustness trends are observed for CREMA-D and Urban Sound. On CREMA-D, at E.R.=0.5, FedMLAC outperforms FedAvg by 8.49%, and FedOPT by over 7%, despite all methods degrading due to noise. On Urban Sound, where non-IID and limited client data exacerbate label noise effects, FedMLAC’s drop remains the lowest at 10.63%, compared to 18.19% and 16.87% for FedAvg and FedProx respectively. This suggests that FedMLAC’s label perturbation-aware components and hybrid aggregation allow it to generalize more robustly under noisy supervision, thereby affirming its practical utility in noisy FL environments.
7 Conclusion and Future Work
This work presents FedMLAC, a unified federated audio classification framework designed to jointly address three central challenges in real-world federated learning (FL): data heterogeneity, model heterogeneity, and data poisoning. By introducing a lightweight, globally shared Plug-in model and a mutual learning mechanism between local and global models, FedMLAC decouples personalization from aggregation, facilitating effective cross-client knowledge transfer while supporting architectural flexibility. Moreover, to counteract the destabilizing effects of noisy or adversarial updates, a novel Layer-wise Pruning Aggregation (LPA) strategy is proposed, enhancing the resilience and stability of the global Plug-in model. Empirical evaluations across four diverse audio benchmarks confirm that FedMLAC consistently outperforms existing FL baselines under clean and noisy conditions, both in homogeneous and heterogeneous model settings.
Despite these advancements, several open challenges remain before practical deployment. Ensuring efficient personalization in the presence of extreme client variability, minimizing communication overhead for low-resource edge devices, and maintaining robustness under adversarial manipulation or privacy-preserving constraints (e.g., differential privacy, secure aggregation) require further exploration. Addressing these general limitations will be critical for building scalable and trustworthy FL systems for audio applications in increasingly complex, decentralized environments.